y-5=7/3(x+4) is in point slope form
A. Write the coordinates of the point
B. Transform the ax+by=c form, where A,B, and C are integer constants
C. Find the x-intercept

Answer by TimothyLamb(4379)

point: (-4, 5)
---
standard form: 7x - 3y = -43
ax + by = c
y - 5 = 7/3(x + 4)
y - 5 = (7/3)x + 28/3
y = (7/3)x + 28/3 + 5
y = (7/3)x + 28/3 + 15/3
y = (7/3)x + 43/3
(7/3)x - y + 43/3 = 0
(7/3)x - (3/3)y + 43/3 = 0
7x - 3y + 43 = 0(3)
7x - 3y + 43 = 0
7x - 3y = -43
---
x-intercept: (-6.14285714, 0)
7x - 3(0) = -43
7x = -43
x = -43/7
y = 0
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php